# Machine learning many-body potentials for charged colloids in primitive 1:1 electrolytes

### v1

@@ III Colloidal System in Low-Polar Solvent
[…]
@@
[…]
@@
[…]
@@
In the main graph of Fig. 4, we compare the effective ML pair potential $U^{ML}_{2}(R)$ of Eq. (14), shown in blue, with the PM potential of Eq. (16), represented by the yellow crosses. It is important to emphasize that the effective ML pair potential is trained on systems with many colloids at much higher packing fractions ($\eta\in[0.001,0.45]$) than the packing fraction used to evaluate the PM potential ($\eta=0.001$) using only 2 colloids. As a result, the effective ML pair potential may incorporate many-body effects, which are strictly absent in the PM potential. Despite this, it remains informative to compare the two potentials. From Fig. 4, we observe that the two potentials agree qualitatively, both representing purely pairwise repulsive interactions; however, the effective ML pair potential exhibits a significantly shorter range than the PM potential. The reason can be traced back to the highest packing fractions included in the ML training, where the ratio $y\equiv ZN/(Vc_{s})$ between the counterion concentration $ZN/V$ and the salt reservoir concentration $c_{s}$ exceeds unity. A simple estimate shows that $y=24\eta(Z\lambda_{B}/\sigma)/(\kappa\sigma)^{2}\simeq 4.5\eta$ for the system parameters considered here. In the counterion-dominated high-$\eta$ regime ($y>1$ so say $\eta>0.22$) screening becomes substantially more efficient compared to the low-$\eta$ regime, where $y<1$. For this reason, we also retrained $U^{ML}(\{{\bf R}\})$ using only configurations with packing fractions restricted to $\eta\in[0.001,0.1]$ such that $y<1$ by a reasonably safe margin. Note that the packing fraction range is still higher than the packing fraction used to evaluate the PM potential. The resulting low-$\eta$ forces are again compared with the PM forces in Fig. 3(b). The resulting ML potential, presented as the orange curve in Fig. 4, more closely approaches the PM pair potential under these low-density conditions. Moreover, the corresponding three-body ML potential in a triangular configuration, shown in the inset, approaches the effective PM potential, with both being purely attractive and exhibiting comparable strength and range. Interestingly, the range of $U_{3}^{ML}(R)$ is again slightly shorter than that obtained from the PM simulations. This suggests that the ML training is overly influenced by the ion-averaged PM forces sampled near the upper bound of the dilute regime at $\eta=0.1$. It is also important to note that small numerical uncertainties in $U^{ML}_{2}$ can lead to relatively large contributions to the effective three-body ML potential $U^{ML}_{3}$.
@@
Next, we go beyond the properties of the two- and three-colloid systems and consider the colloid-colloid radial distribution function $g(R)$ as a function of the center-to-center distance $R$ between the colloidal particles for a variety of packing fractions. To this end, we perform molecular dynamics simulations of an effective colloids-only system using the ML potentials, and measure the radial distribution function. In Fig. 5, we plot $g(R)$ as obtained from primitive model simulations and from molecular dynamics simulations using the ML potential, in (a) trained on all configurations and in (b) only on the configurations in the dilute regime $\eta\in[0.001,0.1]$ as discussed above. Figure 5(a) reveals excellent agreement at the highest three packing fractions $\eta\in[0.1,0.3]$; however, for the dilute system at $\eta=0.01$ the ML prediction is seen to have a much shorter range than the primitive model benchmark. Figure 5(b) shows agreement for $\eta=0.1$, comparable to the agreement observed in panel (a) for the full training set, and for $\eta=0.05$, the $g(r)$’s match well, although not perfectly. For $\eta=0.01$, the agreement improves significantly compared to panel (a) due to the enhanced range of the repulsion, predicted by this version of the ML potential. Nevertheless, a significant deviation from the primitive model structure persists and perfect agreement would probably require training on an even smaller set of low $\eta$.
@@
Figure 5: Radial distribution function $g(R)$ for the colloidal dispersion described in Section III at several packing fractions $\eta$ (see labels) as obtained from primitive model (PM) simulations (lines) and from coarse-grained simulations using ML potentials (triangles), in (a) trained on configurations with packing fractions in the range $\eta\in[0.001,0.45]$ and in (b) on dilute configurations with $\eta\in[0.001,0.1]$.
@@
We conclude that the effective ML potential for this low-salt suspension, where the Debye length of the electrolyte reservoir is of the order of the particle diameter ($\kappa^{-1}=1.54\sigma$ in this case), can effectively capture the ion-averaged forces ${\bf F}^{PM}_{i}$ and the effective colloids-only potential as learned from primitive model simulations. Even though the ML potential only depends on a particle’s local environment, and does not include the long-range interactions of the PM model, it can be used to simulate a system of charged colloids, owing to charge screening in the electrolyte. However, the density dependence remains challenging to capture across the full range of packing fractions. In particular, we find that the high-$\eta$ regime, dominated by counterions, and the low-$\eta$ regime, dominated by the background electrolyte, require separate treatments. Further work is needed to quantify and fully elucidate these behaviors.
@@ IV Salt-Free Asymmetric Electrolyte Potential
In Section III, the investigation targeted an experimental colloidal system falling into the Poisson-Boltzmann regime, which can thus be described accurately by DLVO theory. We now turn our attention to a salt-free colloidal system at low effective temperature $\sigma/\lambda_{B}$ in the strong-coupling regime. This regime is characterised by strong electrostatic interactions and a quasi-two-dimensional layer of counterions condensing on the surface of the colloids rather than the diffuse three-dimensional electric double layer around the colloid, characteristic of the Poisson-Boltzmann regime. The salt-free colloidal dispersion, which can also be seen as an extremely asymmetric electrolyte, consists of colloids with charge valency $Z=50$, and since the system is salt-free with reservoir screening parameter $\kappa\sigma=0$, we add $50$ charge-neutralizing counterions for every colloid, with ion-to-colloid size ratio $\sigma_{i}/\sigma=0.05$. Since we are interested in the strongly-coupled electrostatic regime, we simulate at low temperatures in the range $\sigma/\lambda_{B}\in[1.2,2.4]$.
@@
Figure 6: (a) The potential of mean force $U^{PM}_{2}(R)$ (crosses) and the potential $U^{ML}_{2}(R)$ predicted by our ML model (solid lines), for a system consisting of colloids with charge valency $Z=50$ and counterions, with ion-to-colloid size ratio $\sigma_{i}/\sigma=0.05$, and effective temperatures $\sigma/\lambda_{B}\in[1.2,2.4]$. The ML potential is trained on primitive model simulation of $N=32$ colloids with packing fractions in the range $\eta\in[0.001,0.45]$ and may incorporate many-body effects, whereas the PM potential is strictly a two-body potential albeit density-dependent in this salt-free case. (b) Radial distribution function $g(R)$ for the salt-free colloidal dispersion described above, at packing fraction $\eta=0,2$ as obtained from primitive model simulations (lines) and from coarse-grained simulations using ML potentials (triangles)
@@
In Fig. 6(a), we compare the effective ML pair potential $U_{2}^{ML}(R)$ (solid lines) with the potential of mean force $U_{2}^{PM}(R)$ (crosses) obtained from primitive model simulations for various values of $\sigma/\lambda_{B}$. The potential of mean force $U_{2}^{PM}(R)$ is determined using primitive model simulations with two colloids and $100$ counterions in a cubic simulation box with side length $L=6\sigma$, corresponding to a colloid packing fraction $\eta=0.0048$. In these simulations, the colloids are held fixed at different separations while the forces acting on them are thermally averaged over the counterion configurations. The PM potential $U_{2}^{PM}(R)$ is then calculated using Eq. (16). From both potentials in Fig. 6(a), we observe features characteristic of interactions in the strong coupling regime. At high temperatures, $\sigma/\lambda_{B}=2.4$, the two-body potentials are purely repulsive. As the temperature decreases, an attractive well starts to form, until this well becomes the dominant feature at $\sigma/\lambda_{B}=1.2$. Surprisingly, the PM potential and the ML potential agree qualitatively at $\sigma/\lambda_{B}=1.2$, since the attractive well emerges at the same temperature and with comparable prominence for both potentials. In this regime, the counterions are condensed to the surface of the colloids, which explains the perfect match between the effective ML potential and the PM potential. However, quantitative differences remain at higher values of $\sigma/\lambda_{B}$. This is expected, as the effective ML potentials may include many-body effects as they are trained on systems with many colloids at relatively high densities. Furthermore, within the salt-free case, the potential of mean force does not have a well-defined dilute limit and remains density-dependent upon dilution. Fig. 6 shows that the PM potential exhibits a pronounced long-range repulsion, whereas the ML potential shows this feature only weakly. The ML potential does not exhibit these long-range features because a relatively high weight is assigned to the high-$\eta$ regime, where short-range interactions and large forces dominate. However, the ML potential effectively captures the changes in behavior associated with varying temperature.
@@
Figure 7: (a) The effective cation-cation pair potential $U_{2}(R)$ as predicted by our ML models (blue solid lines), according to the Debye-Hückel Theory $U^{DH}=\lambda_{B}\exp{[-\kappa R]}/R$ (red solid line), and measured in primitive model (PM) simulations (yellow crosses) for a 1:1 electrolyte at effective temperature $\sigma_{i}/\lambda_{B}=0.5$ and $\eta=0.01$. The ML potential is trained on configurations with packing fractions in the range $\eta\in[0.001,0.1]$. (b) Cation-cation radial distribution function $g_{++}(R)$ in an electrolyte at several packing fractions $\eta$ (see labels) as obtained from primitive model simulations (lines) and from coarse-grained simulations using ML potential (triangles).
@@ VI Conclusion
In this work, we have applied a machine-learning framework to generate potentials that mimic effective interactions of charged particles, either colloids or ions, suspended in an electrolyte. We considered colloid-colloid, colloid-ion, and ion-ion interactions within the primitive model, where colloidal and ionic particles are treated explicitly, and the solvent is represented as a dielectric continuum. Effective like-charge attractions between colloidal particles can arise in primitive model simulations when the colloidal surface charge density is high and the effective temperature is low, indicative of the strongly coupled electrostatic regime. However, large-scale primitive model simulations of colloidal dispersions require the explicit treatment of a large number of ions and the evaluation of long-range electrostatic interactions, making these simulations computationally demanding. This motivated us to generate machine-learned potentials that include many-body components, enabling us to perform signficantly faster simulations while retaining the essential physics of the system.
@@
The ML framework developed here to accelerate simulations of charged particles are useful for predicting the phase behavior of dispersions of highly charged colloids, even in regimes of low or zero salt. However, several aspects of how salt affects phase behavior remains to be explored, as both experiments Larsen and Grier (1997); Gomez et al. (2009); Wang et al. (2024) and simulations Hynninen and Panagiotopoulos (2009) have shown that finite salt concentrations can change the effective colloid-colloid interactions in unexpected ways. For highly charged colloids, the effect of adding salt can also be studied by training ML potentials across a range of screening lengths, corresponding to various salt concentrations, at physically relevant coupling parameters. By comparing these potentials, it becomes possible to study like-charge attractions within the many-body interactions, analyze their dependence on screening length, and determine whether these attractive terms can drive phase coexistence. This will be the topic of a future publication.

### v2

@@ III Colloidal System in Low-Polar Solvent
Due to the linearity of our ML model and the way it is trained, we also have direct access to the scalar coarse-grained many-body potential $U^{ML}(\{{\bf R}\})$. To gauge the accuracy of the effective ML potential $U^{ML}(\{{\bf R}\})$— trained on the PM-generated force-vector ${\bf f}^{PM}$—we evaluate the effective ML potential for a system consisting of only two colloids $(N=2$) at positions ${\bf R}_{1}$ and ${\bf R}_{2}$, such that we can define the effective ML pair potential
@@
$U_{2}^{PM}(R)=\int_{R}^{\infty}dR^{\prime}\frac{\mathbf{F}^{PM}_{i}(R^{\prime})\cdot\mathbf{R}_{ij}}{R^{\prime}},$ (16)
@@
$U_{3}^{PM}(R)$ $=$ $\frac{2}{\sqrt{3}}\int_{R}^{\infty}dR^{\prime}\frac{\mathbf{F}^{PM}_{i}(R^{\prime})\cdot\mathbf{R}_{ij}+\mathbf{F}^{PM}_{i}(R^{\prime})\cdot\mathbf{R}_{ik}}{R^{\prime}}$ (17)
@@
In the main graph of Fig. 4, we compare the effective ML pair potential $U^{ML}_{2}(R)$ of Eq. (14), shown in blue, with the PM potential of Eq. (16), represented by the yellow crosses. It is important to emphasize that the effective ML pair potential is trained on systems with many colloids at much higher packing fractions ($\eta\in[0.001,0.45]$) than the packing fraction used to evaluate the PM potential ($\eta=0.001$) using only 2 colloids. As a result, the effective ML pair potential may incorporate many-body effects, which are strictly absent in the PM potential. Despite this, it remains informative to compare the two potentials. From Fig. 4, we observe that the two potentials agree qualitatively, both representing purely pairwise repulsive interactions; however, the effective ML pair potential exhibits a significantly shorter range than the PM potential. The reason can be traced back to the highest packing fractions included in the ML training, where the ratio $y\equiv ZN/(Vc_{s})$ between the counterion concentration $ZN/V$ and the salt reservoir concentration $c_{s}$ exceeds unity. A simple estimate shows that $y=24\eta(Z\lambda_{B}/\sigma)/(\kappa\sigma)^{2}\simeq 4.5\eta$ for the system parameters considered here. In the counterion-dominated high-$\eta$ regime ($y>1$ so say $\eta>0.22$) screening becomes substantially more efficient compared to the low-$\eta$ regime, where $y<1$. For this reason, we also retrained $U^{ML}(\{{\bf R}\})$ using only configurations with packing fractions restricted to $\eta\in[0.001,0.1]$ such that $y<1$ by a reasonably safe margin, resulting in a total of 204 configurations. Note that the packing fraction range is still higher than the packing fraction used to evaluate the PM potential. The resulting low-$\eta$ forces are again compared with the PM forces in Fig. 3(b). We find that the ML potential reproduces the PM forces, though less successfully than for the full potential, as indicated by the lower coefficient of determination, $R^{2}=0.865$. The resulting ML potential, presented as the orange curve in Fig. 4, more closely approaches the PM pair potential under these low-density conditions. Moreover, the corresponding three-body ML potential in a triangular configuration, shown in the inset, approaches the effective PM potential, with both being purely attractive and exhibiting comparable strength and range. An attractive three-body contribution to the potential, as observed in simulations Lin et al. (2021) and in experiments,Dobnikar et al. (2004); Brunner et al. (2004); Merrill, Sainis, and Dufresne (2009) can promote particle aggregation, with many-body interactions becoming particularly significant in dense clusters of particles. Interestingly, the range of $U_{3}^{ML}(R)$ is again slightly shorter than that obtained from the PM simulations. This suggests that the ML training is overly influenced by the ion-averaged PM forces sampled near the upper bound of the dilute regime at $\eta=0.1$. It is also important to note that small numerical uncertainties in $U^{ML}_{2}$ can lead to relatively large contributions to the effective three-body ML potential $U^{ML}_{3}$.
@@
Next, we go beyond the properties of the two- and three-colloid systems and consider the colloid-colloid radial distribution function $g(R)$ as a function of the center-to-center distance $R$ between the colloidal particles for a variety of packing fractions. To this end, we perform molecular dynamics simulations of an effective colloids-only system using the ML potentials, and measure the radial distribution function. In Fig. 5, we plot $g(R)$ as obtained from primitive model simulations and from molecular dynamics simulations using the ML potential, in (a) trained on all configurations and in (b) only on the configurations in the dilute regime $\eta\in[0.001,0.1]$ as discussed above. Figure 5(a) reveals excellent agreement at the highest three packing fractions $\eta\in[0.1,0.3]$. However, the ML potential fails to reproduce the primitive model simulation results at low densities; for the dilute system at $\eta=0.01$ the ML prediction exhibits a much shorter interaction range than the primitive model benchmark. Figure 5(b) shows agreement for $\eta=0.1$, comparable to the agreement observed in panel (a) for the full training set, and for $\eta=0.05$, the $g(r)$’s match well, although not perfectly. For $\eta=0.01$, the agreement improves significantly compared to panel (a) due to the enhanced range of the repulsion, predicted by this version of the ML potential. Nevertheless, a significant deviation from the primitive model structure persists and perfect agreement would probably require training on an even smaller set of low $\eta$.
@@
Figure 5: Radial distribution function $g(R)$ for the colloidal dispersion described in Section III at several packing fractions $\eta$ (see labels) as obtained from primitive model (PM) simulations (lines) and from coarse-grained simulations using ML potentials (triangles), in (a) trained on configurations with packing fractions in the range $\eta\in[0.001,0.45]$, with the curves shifted upwards for $\eta=0.1$, $\eta=0.2$, and $\eta=0.3$ by +1+1, +2+2, and +3+3, respectively, and in (b) on dilute configurations with $\eta\in[0.001,0.1]$, with the curves shifted upwards for $\eta=0.05$ and $\eta=0.1$ by +1+1 and +2+2, respectively
@@
We conclude that the effective ML potential for this low-salt suspension, where the Debye length of the electrolyte reservoir is on the order of the particle diameter ($\kappa^{-1}=1.54\sigma$ in this case), can effectively capture the ion-averaged forces ${\bf F}^{PM}_{i}$ and the effective colloids-only potential as learned from primitive model simulations for relatively dense colloid suspensions, although its performance becomes worse at low densities. Even though the ML potential only depends on a particle’s local environment, and does not include the long-range interactions of the PM model, it can be used to simulate a system of charged colloids, owing to charge screening in the electrolyte. However, the density dependence remains challenging to capture across the full range of packing fractions. In particular, we find that the high-$\eta$ regime, dominated by counterions, and the low-$\eta$ regime, dominated by the background electrolyte, require separate treatments. Further work is needed to quantify and fully elucidate these behaviors. The density dependence of the potential should be explicitly included during ML potential training. This can be achieved by incorporating density-dependent weights in the linear regression, using neural networks trained also on density, or by including higher-body descriptors,Nguyen and Sema (2024) long-range descriptors,Loche et al. (2025) or Ewald summation techniques.Kim et al. (2025)
@@ IV Salt-Free Asymmetric Electrolyte Potential
In Section III, the investigation targeted an experimental colloidal system falling into the Poisson-Boltzmann regime, which can thus be described accurately by DLVO theory. We now turn our attention to a salt-free colloidal system at low effective temperature $\sigma/\lambda_{B}$ in the strong-coupling regime. This regime is characterised by strong electrostatic interactions and a quasi-two-dimensional layer of counterions condensing on the surface of the colloids rather than the diffuse three-dimensional electric double layer around the colloid, characteristic of the Poisson-Boltzmann regime. The salt-free colloidal dispersion, which can also be seen as an extremely asymmetric electrolyte, consists of colloids with charge valency $Z=50$, and since the system is salt-free with reservoir screening parameter $\kappa\sigma=0$, we add $50$ charge-neutralizing counterions for every colloid, with ion-to-colloid size ratio $\sigma_{i}/\sigma=0.05$. Since we focus on the strongly coupled electrostatic regime, we perform simulations at low temperatures and generate four separate ML potentials corresponding to $\sigma/\lambda_{B}=1.2$, $\sigma/\lambda_{B}=1.6$, $\sigma/\lambda_{B}=2.0$, and $\sigma/\lambda_{B}=2.4$.
@@
Figure 6: (a) The potential of mean force $U^{PM}_{2}(R)$ (crosses) and the potential $U^{ML}_{2}(R)$ predicted by our ML model (solid lines), for a system consisting of colloids with charge valency $Z=50$ and counterions, with ion-to-colloid size ratio $\sigma_{i}/\sigma=0.05$, and effective temperatures $\sigma/\lambda_{B}\in[1.2,2.4]$. The ML potential is trained on primitive model simulation of $N=32$ colloids with packing fractions in the range $\eta\in[0.001,0.45]$ and may incorporate many-body effects, whereas the PM potential is strictly a two-body potential, since it was determined using simulations of two isolated colloids, albeit density-dependent in this salt-free case. (b) Radial distribution function $g(R)$ for the salt-free colloidal dispersion described above, at packing fraction $\eta=0.2$ as obtained from primitive model simulations (lines) and from coarse-grained simulations using ML potentials (triangles). The curves for $\sigma/\lambda_{B}=2.0$, $\sigma/\lambda_{B}=1.6$, $\sigma/\lambda_{B}=1.2$ are shifted upwards by +2+2, +4+4, and +6+6, respectively.
@@
In Fig. 6(a), we compare the effective ML pair potential $U_{2}^{ML}(R)$ (solid lines) with the potential of mean force $U_{2}^{PM}(R)$ (crosses) obtained from primitive model simulations for various values of $\sigma/\lambda_{B}$. The potential of mean force $U_{2}^{PM}(R)$ is determined using primitive model simulations with two colloids and $100$ counterions in a cubic simulation box with side length $L=6\sigma$, corresponding to a colloid packing fraction $\eta=0.0048$. In these simulations, the colloids are held fixed at different separations while the forces acting on them are thermally averaged over the counterion configurations. The PM potential $U_{2}^{PM}(R)$ is then calculated using Eq. (16). From both potentials in Fig. 6(a), we observe features characteristic of interactions in the strong coupling regime. At high temperatures, $\sigma/\lambda_{B}=2.4$, the two-body potentials are purely repulsive. As the temperature decreases, an attractive well starts to form, until this well becomes the dominant feature at $\sigma/\lambda_{B}=1.2$. Surprisingly, the PM potential and the ML potential show qualitative agreement at $\sigma/\lambda_{B}=1.2$, as the attractive well appears at the same temperature and with comparable depth in both cases. In this regime, the counterions are condensed to the surface of the colloids, which explains the reasonable match between the effective ML potential and the PM potential. However, quantitative differences develop at higher values of $\sigma/\lambda_{B}$ — as expected since in the salt-free case the potential of mean force only reaches its dilute limit—bare colloid-colloid Coulomb interactions—at extremely low densities. In fact we estimate on the basis of Ref.57 that the potential of mean force remains density-dependent upon dilution down to $\eta\simeq 10^{-15}$. Furthermore, we note that the qualitative differences between ML and PM potentials may also arise from many-body effects that are included in the effective ML potentials as they are trained on systems with many colloids at relatively high densities and thus high counterion concentrations. However, the ML potential does capture the qualitative change in colloid interactions with decreasing temperature, showing a reduction in repulsion and the emergence of a short-ranged attractive well.
@@
Figure 7: (a) The effective cation-cation pair potential $U_{2}(R)$ as predicted by our ML models (blue solid lines), according to the Debye-Hückel Theory $U^{DH}=\lambda_{B}\exp{[-\kappa R]}/R$ (red solid line), and measured in primitive model (PM) simulations (yellow crosses) for a 1:1 electrolyte at effective temperature $\sigma_{i}/\lambda_{B}=0.5$ and $\eta=0.01$. The ML potential is trained on configurations with packing fractions in the range $\eta\in[0.001,0.1]$. (b) Cation-cation radial distribution function $g_{++}(R)$ in an electrolyte at several packing fractions $\eta$ (see labels) as obtained from primitive model simulations (lines) and from coarse-grained simulations using ML potential (triangles). The curves for $\eta=0.01$ and $\eta=0.05$ are shifted upwards by +1+1 and +2+2, respectively.
@@ VI Conclusion
In this work, we have applied a machine-learning framework to generate potentials that mimic effective interactions of charged particles, either colloids or ions, suspended in an electrolyte. We considered colloid-colloid, colloid-ion, and ion-ion interactions within the primitive model, where colloidal and ionic particles are treated explicitly, and the solvent is represented as a dielectric continuum. Effective like-charge attractions between colloidal particles can arise in primitive model simulations when the colloidal surface charge density is high and the effective temperature is low, indicative of the strongly coupled electrostatic regime. However, large-scale primitive model simulations of colloidal dispersions require the explicit treatment of a large number of ions and the evaluation of long-range electrostatic interactions, making these simulations computationally demanding. This motivated us to generate machine-learned potentials that include many-body components, enabling us to perform significantly faster simulations while retaining the essential physics of the system.
@@
The ML framework developed here enables large-scale simulations of charged particles, allowing us to probe the phase behavior of highly charged colloidal dispersions, even at low or zero salt. However, the current approach is limited by the need to generate a separate ML potential for each temperature and by its poor representation of low colloid densities. These limitations could be addressed in future work by incorporating temperature- and density-dependent weights in the linear regression, or by using more advanced techniques such as neural networks, higher-body descriptors,Nguyen and Sema (2024) long-range descriptors,Loche et al. (2025) or Ewald summation methods.Kim et al. (2025) Nevertheless, the framework is still instrumental for exploring several aspects of how salt affects phase behavior, as both experiments Larsen and Grier (1997); Gomez et al. (2009); Wang et al. (2024) and simulations Hynninen and Panagiotopoulos (2009) have shown that finite salt concentrations can change the effective colloid-colloid interactions in unexpected ways. For highly charged colloids, the effect of adding salt can also be studied by training ML potentials across a range of screening lengths, corresponding to various salt concentrations, at physically relevant coupling parameters. By comparing these potentials, it becomes possible to study like-charge attractions within the many-body interactions, analyze their dependence on screening length, and determine whether these attractive terms can drive phase coexistence. In future work, we will examine how attractive three-body interactions and added salt influence the phase behavior of charged colloids. Finally, we note that interfaces between coexisting phases are as intriguing as they are challenging, since the density-dependent two-body repulsions and three-body attractions identified for certain parameter regimes were derived from training on quasi-homogeneous rather than strongly heterogeneous states. Assessing to what extent such density-dependent ML potentials can capture interfacial phenomena in phase-separated colloidal dispersions remains an open and challenging problem.